# WanJuan-CC: A Safe and High-Quality Open-sourced English Webtext Dataset

### 29 Feb 2024 v1

@@ 3.2 Data Extraction
Each Common Crawl dump contains the original content of the crawled web pages, such as HTML, CSS, JS, and URLs of multimedia resources. To obtain relatively clean and coherent training corpora, we need to extract text information from these raw data. Additionally, at this stage, we can also perform language classification. We employed the trafilacture method[18] for text extraction. Subsequently, we utilized the pycld2 library to classify the language of the extracted text. As this study is solely concerned with English data, we retained only the data marked as “en”, discarding data in other languages.
@@ 3.6.2 Fluency Classification
3.7 Data Quality Assessment and Feedback Loop
After processing with the aforementioned data processing pipeline, we obtain a High-Quality Data. The design of rules and selection of filtering thresholds are manually evaluated and adjusted to achieve the desired effect. To ensure data quality, we have implemented a comprehensive data quality assessment and feedback mechanism in our engineering practice. It is widely recognized that assessing the quality of the Common Crawl dataset presents significant challenges, primarily due to its large volume, diverse problem types, vague quality definitions, high degree of subjectivity, and scattered features.
@@ 3.7 Data Quality Assessment and Feedback Loop
  • Quality Signal: Format
    • Ratio of data items in the dataset that contain errors in formatting (such as formulas, tables, markdown).
    • Ratio of data items in the dataset that contain special characters (such as spaces, line breaks, #, invisible characters).
@@
  • Quality Signal: Consistency
@@
  • Quality Signal: Unsafety
@@
In order to conduct a comprehensive, accurate, and efficient data quality assessment, we conceptualize the process into three distinct stages: metric identification and definition, computation methods for these metrics, and determination of metric thresholds. To address these stages, we employ three disparate quality assessment methodologies: 1) Human Evaluation Method; 2) Quality Signal Evaluation Method; 3) Model Evaluation Method.
@@
Quality Signal Evaluation Method: To accurately and efficiently gauge the extent of data quality issues within the entire Common Crawl dataset, we have devised machine rules that are independent of the data cleaning stage. These rules automatically evaluate the proportion of poor-quality samples in the entire dataset based on the quality indicators identified through the Human Evaluation Method. We define this proportion as $P=\frac{A}{B}$, where $A$ represents the number of poor-quality samples, and $B$ denotes the total number of samples in the dataset.
However, not all quality indicators can yield an accurate proportion of the entire dataset via machine rules, especially for complex issues such as texts containing leading or trailing promotional information. For these types of quality indicators, we estimate the proportion of poor-quality samples as $\tilde{P}=\frac{A^{\prime}}{B^{\prime}}$, where $A^{\prime}$ is the number of poor-quality samples identified by Human Evaluation, and $B^{\prime}$ is the total number of sampled instances.
Model Evaluation Method: We have obtained the proportions of quality indicators through both human and machine methods, which serve to measure the severity of issues. However, determining what proportion is deemed serious or not requires an independent threshold for each quality indicator. We initially set these thresholds based on our experience, then update the signals and thresholds in response to feedback from downstream Model Evaluation. In the Model Evaluation phase, we have designed a test dataset comprising approximately 600 prompts, spanning over 20 types in both Chinese and English. We then conduct a comprehensive assessment of the model-generated content using both subjective evaluation and Quality Signal Evaluation. This process allows us to explore the impact of data quality issues on model performance, thereby facilitating updates to the quality indicators and thresholds within our quality standards.

### 01 Mar 2024 v2

@@ 3.2 Data Extraction
Each Common Crawl dump contains the original content of the crawled web pages, such as HTML, CSS, JS, and URLs of multimedia resources. To obtain relatively clean and coherent training corpora, we need to extract text information from these raw data. Additionally, at this stage, we can also perform language classification. We employed the Trafilatura method[18] for text extraction. Subsequently, we utilized the pycld2 library to classify the language of the extracted text. As this study is solely concerned with English data, we retained only the data marked as “en”, discarding data in other languages.
@@ 3.6.2 Fluency Classification
3.7 Data Quality Evaluation
After processing with the aforementioned data processing pipeline, we obtain a High-Quality Data. The design of rules and selection of filtering thresholds are manually evaluated and adjusted to achieve the desired effect. To ensure data quality, we have implemented a comprehensive data quality evaluation and feedback mechanism in our engineering practice. It is widely recognized that assessing the quality of the Common Crawl dataset presents significant challenges, primarily due to its large volume, diverse problem types, vague quality definitions, high degree of subjectivity, and scattered features.
@@ 3.7 Data Quality Evaluation
  • Quality Signal: Understandability
    • Ratio of data items in the dataset that contain errors in formatting (such as formulas, tables, markdown, etc.).
    • Ratio of data items in the dataset that contain special characters (such as high proportion of spaces, line breaks, #, invisible characters).
@@
  • Quality Signal: Relevance
@@
  • Quality Signal: Security
@@
In order to conduct a comprehensive, accurate, and efficient data quality evaluation, we conceptualize the process into three distinct stages: metric identification and definition, computation methods for these metrics, and determination of metric thresholds. To address these stages, we employ three disparate quality assessment methodologies: 1) Human Evaluation Method; 2) Quality Signal Evaluation Method; 3) Model Evaluation Method.
@@
Quality Signal Evaluation Method: To accurately and efficiently gauge the extent of data quality issues within the entire Common Crawl dataset, we have devised machine rules that are independent of the data cleaning stage. These rules automatically evaluate the proportion of poor-quality samples in the entire dataset based on the quality indicators identified through the Human Evaluation Method. We define this proportion as $X=\frac{A}{B}$, where $A$ represents the number of poor-quality samples, and $B$ denotes the total number of samples in the dataset.
However, not all quality indicators can yield an accurate proportion of the entire dataset via machine rules, especially for complex issues such as texts containing leading or trailing promotional information. For these types of quality indicators, we estimate the proportion of poor-quality samples as $\tilde{X}=\frac{A^{\prime}}{B^{\prime}}$, where $A^{\prime}$ is the number of poor-quality samples identified by Human Evaluation, and $B^{\prime}$ is the total number of sampled instances.
Model Evaluation Method: We have obtained the proportions of quality indicators through both human and machine methods, which serve to measure the severity of issues. However, determining what proportion is deemed serious or not requires an independent threshold for each quality indicator. We initially set these thresholds based on our experience, then update the signals and thresholds in response to feedback from downstream Model Evaluation. Firstly, we will use the 1B and 3B experimental models trained in Section 4.4 for evaluation. In the Model Evaluation phase, we have designed a test dataset comprising approximately 600 prompts, spanning over 20 types in both Chinese and English. We then conduct a comprehensive assessment of the model-generated content using both subjective evaluation and Quality Signal Evaluation. This process allows us to explore the impact of data quality issues on model performance, thereby facilitating updates to the quality indicators and thresholds within our quality standards.